What is the speed of sound, if air temperature is -20 Β°C?
1
Expert's answer
2021-10-20T15:58:05-0400

Speed of sound in air is directly proportional to square root of absolute temperature.

The speed of sound in air at 27 degress is 340m/s.

"T_1 = 27 + 273 =300 \;K \\

T_2 = -20 +273 = 253 \;K \\

v_2 = v_1 \times \sqrt{\frac{T_2}{T_1}} \\

v_2 = 340 \times \sqrt{\frac{253}{300}} \\

v_2 =312.23 \; m/s"